[ x-1 ] + [ x-5 ] = 20

Grade:11

3 Answers

Shrey
10 years ago
x-1+x-5=20 2x-6=20 2x=20+6 2x=26 x=26/2 x=13 Please approve my answer by clicking approved if you liked it.
shivam chopra
10 years ago
[ x-1 ] + [ x-5 ] = 20 . " [ x-1 ].. " these brackets are of greatest integer function!! if it was that much simpler then why should i`ve asked it , for querry see chapter functions of class 12th
vaanya
10 years ago
if u equate each bracket to 0 then u get x=1,5 which would give you 3 cases: x>=5 , 1<=x<5 , x<=1 (u can try it out by putting a value or u can remember this as a rule...when the case gives a positive value inside bracket you can simply take out the integer with the sign without worrying abt GIF...but if the case gives a negative value inside GIF bracket then u take out the integer from the bracket with the sign AND ALSO subtract 1 for the bracket) Case 1: x>=5, (this gives positive values for both brackets) [x-1]+[x-5]=20 [x]-1+[x]-5=20 [x]=13 x belongs to [13,14) Case 2: 1<=x<5 (this gives a positive value for the 1t bracket and a negative value for the second bracket) [x-1]+[x-5]=20 [x]-1+[x]-5`-1`=20 (note the part in apostrophes) [x]=27/2 which is not an integer and GIF can never be a non-integer so this case gives us no value. Case 3: x<1 (this gives both brackets as negative) [x-1]+[x-5]=20 [x]-1`-1`+[x]-5`-1`=20 [x]=14 x belongs to [14,15) so the complete answer is [13,15) i hope u got what i am trying to say. approve if u understood and liked the answer.